# Improving Efficiency of Sampling-based Motion Planning via Message-Passing Monte Carlo

### v1

@@ Abstract
I INTRODUCTION
@@
Instead of refining motion planning algorithms, our work targets the fundamental building block underlying these approaches—the core sampling process itself. We introduce a novel sampling strategy based on Message-Passing Monte Carlo (MPMC) [11], a graph neural network architecture trained to generate low discrepancy point sets on unit hypercubes of arbitrary dimension. We expand the generic MPMC algorithm with the introduction of a novel training objective tailored to high-dimensional spaces, ensuring that generated points are optimally distributed and scalable.
This leads to a potent unbiased state sampling technique, that can be seamlessly integrated into any sampling-based planner, and that requires neither conditioning on the workspace description, nor a steer function, nor past examples, nor start and goal information. This approach provides strong theoretical guarantees and it also outperforms traditional techniques across various benchmarks, in environments of varying complexity and dimensionality.
Our contributions are stated as follows:
@@
Our MPMC-based sampling strategy offers a powerful and versatile alternative to traditional sampling methods. By targeting the sampling process directly, we open up new possibilities for improving the efficiency of a wide array of motion planning algorithms in both simulated and real-world robotic systems.
II Related work
II-A Learning for sampling-based planners
@@
II-B Low-discrepancy constructions
Over the past century, numerous constructions of low-discrepancy point sets and sequences have been proposed. Most of the early constructions are deeply rooted in number theory and abstract algebra. A widely used building block of previous low-discrepancy constructions is the one-dimensional van der Corput sequence [18] in base $b$, which is generalized to a higher dimensional setting via the Halton sequence [19]. Each of the $d$ coordinates in a Halton sequence corresponds to a distinct van der Corput sequence in base $b$, with the bases selected to be co-prime. Faure sequences, as introduced in [20], offer a similar construction to Halton sequences but incorporate permutations of the digits in base $b$.
@@ II-B Low-discrepancy constructions
Several approaches have recently emerged targeting the construction of points for fixed dimension and number of points. In [28], new low-discrepancy point sets were proposed by optimizing permutations applied to the Halton sequence. Another approach, known as subset selection, was introduced to select $k<N$ points from an $N$-element set that minimize the discrepancy. An exact algorithm for this selection was presented in [29], while a swap-based heuristic approach was employed in [30]. Additionally, [31] proposed a non-linear programming method to generate point sets with optimal star-discrepancy for fixed dimension and number of points. However, this approach faces significant computational challenges, limiting its practical application to finding optimal sets for only up to $21$ points in two dimensions and $8$ points in three dimensions. Message-Passing Monte Carlo (MPMC) falls into this class of constructions by optimizing the discrepancy of point sets that are fixed for a chosen dimension and number of points. However, MPMC significantly differs from all previous methods by its explicit use of machine learning. Moreover, it has been shown in [11] that MPMC generates point sets with significantly better distributional properties compared to any previous method, reaching optimal or near-optimal discrepancy.
III METHODS
@@
III-A Message-Passing Monte Carlo Sampling
Message-Passing Monte Carlo (MPMC) [11] leverages Graph Neural Networks (GNNs) to generate point sets that cover the space in a uniform manner. This uniformity can be assessed through measures of irregularity termed discrepancy. While there exist a plethora of different uniformity measures, we focus on the ${\mathcal{L}}_{p}$-discrepancy here. That is, given a set of points $\{{\bf X}_{i}\}_{i=1}^{N}$ in the unit hypercube $[0,1]^{d}$ and $p\geq 1$, the ${\mathcal{L}}_{p}$-discrepancy is defined as,
${\mathcal{L}}_{p}^{p}(\{{\bf X}_{i}\}_{i=1}^{N}):=$ (1)
@@ III-A Message-Passing Monte Carlo Sampling
where #($\#(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,{\bf x}))$ counts how many points of $\{{\bf X}_{i}\}_{i=1}^{N}$ fall inside the box $[0,{\bf x})=\prod_{i=1}^{d}[0,x_{i})$ for ${\bf x}=(x_{1},\ldots,x_{d})\in[0,1]^{d}$, and $\mu(\cdot)$ denotes the usual Lebesgue measure. A differentiable closed-form solution to the high-dimensional integral in (LABEL:eq:L2_disc) for the case of $p=2$, known as Warnock’s formula [32], can be computed in $\mathcal{O}(N^{2}d)$. Consequently, MPMC utilizes the ${\mathcal{L}}_{2}$-discrepancy as the objective function minimized during training to generate point sets with low discrepancy. The full MPMC model transforming random input points into low-discrepancy point sets via deep GNNs is depicted in Fig. 1.
III-A1 Fast MPMC training in higher dimensions
As described in [11], the ${\mathcal{L}}_{2}$-discrepancy fails to distinguish random from highly uniform points in high dimensions for moderate amount of points. Therefore, it was suggested in [11] to minimize the Hickernell ${\mathcal{L}}_{2}$-discrepancy $D_{H,2}$ [33] in this case, that sums over all ${\mathcal{L}}_{2}$-discrepancies of $k$-dimensional projections, with $1\leq k\leq d$. A straightforward implementation of $D_{H,2}$ has a computational complexity of $\mathcal{O}((2^{d}-1)N^{2}d)$, which would involve evaluating the $\mathcal{L}_{2}$-discrepancy for over a billion projections for $d=32$, a relatively small number of points. To mitigate this issue, the authors in [11] proposed limiting the summation to a smaller, randomly selected subset of projections. In contrast, in this work we leverage the closed-form solution of the Hickernell ${\mathcal{L}}_{2}$-discrepancy from [34] that can be computed again in $\mathcal{O}(N^{2}d)$ instead of $\mathcal{O}((2^{d}-1)N^{2}d)$. More concretely,
$=\left(\frac{4}{3}\right)^{d}-\frac{2}{N}\sum_{i=1}^{N}\prod_{k=1}^{d}\left(\frac{3}{2}-\frac{{\bf X}_{i,k}^{2}}{2}\right)$
$+\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\prod_{k=1}^{d}[2-\max({\bf X}_{i,k},{\bf X}_{j,k})],$
where ${\bf X}_{i,k}$ is the $k$-th entry of ${\bf X}_{i}$, and $\{{\bf X}^{(s)}_{i}\}_{i=1}^{N}$ is the projection of $\{{\bf X}_{i}\}_{i=1}^{N}$ onto $[0,1]^{|s|}$. This enables very fast training for generating MPMC points in higher dimensions that minimize the discrepancy over all projections.
III-B Motion planning via Probabilistic Road Maps
The algorithm begins by sampling a fixed number $N$ of points from the entire space to form the set $S$. Next, we prune out milestones from the initial set $S$ that do not fall in free space, resulting in a refined set of valid samples $S_{v}$, which is combined with the start and goal nodes $\{{\bf X}_{\text{start}},{\bf X}_{\text{goal}}\}$ to obtain the PRM graph vertex set $V$. For each node $v$ in $V$, the algorithm identifies nearby nodes within a radius $r_{N}$, with the set of such points denoted $P_{\text{near}}$. For each neighbor $p$ in $P_{\text{near}}$, the algorithm checks for collision-free paths between $v$ and $p$. If the path is free of collisions, a bidirectional edge is added between $v$ and $p$ to the edge set $E$. Finally, the algorithm attempts to find a shortest path from ${\bf X}_{\text{start}}$ to ${\bf X}_{\text{goal}}$ and returns the path if one exists; otherwise, it indicates failure.
Algorithm 1 ps-PRM Algorithm
1:$S\leftarrow\texttt{Sample}(N)$
2:$S_{v}\leftarrow\texttt{PruneInvalid}(S)$
3:$V\leftarrow\{{\bf X}_{\text{start}},{\bf X}_{\text{goal}}\}\cup S_{v}$
4:$E\leftarrow\emptyset$
@@ III-B Motion planning via Probabilistic Road Maps
III-C Theoretical guarantee
Next, we outline a theoretical justification of minimizing discrepancy as a means to improving efficiency of sampling-based motion planning. To this end, we introduce another uniformity measure known as dispersion, which is commonly used in assessing the efficiency of sampling-based motion planning. Given a point set $\{{\bf X}^{s}_{i}\}_{i=1}^{N}$ in $[0,1]^{d}$ and $p\geq 1$, the $l_{p}$-dispersion is then defined as,
@@ III-C Theoretical guarantee
with the ${\mathcal{L}}_{\infty}$-discrepancy. Based on this and following [36], one can provide a deterministic sampling guarantee. Namely, using the PRM planning algorithm, assuming some mild assumptions are satisfied (e.g., feasible path $\delta$-clearance), and choosing a radius $r_{N}$ of the PRM algorithm based on $N$ samples according to,
@@
for some $\alpha>1$ and $r_{N}$ satisfying conditions corresponding to the path $\delta$-clearance, one can guarantee that the cost of the returned path is within a factor of $\frac{1}{\alpha-1}$ of the optimal $\delta$-clear path. Clearly, a low discrepancy induces a guarantee of the resulting path to be close to the optimal $\delta$-clear path.
IV Experimental Evaluation
We evaluate our motion planning solution in simulation, using a variety of environments ranging from 2D mazes to higher dimensional spaces and in a physical setting with a UR5 robot arm. We assess performance according to the number of sampled points and success rate in finding a path. Code for generating MPMC points can be found at https://github.com/tk-rusch/MPMC
IV-A Sampling methods
@@
The performance of our method is evaluated against Uniform sampling and also against Halton [37] and Sobol [38] (for the 2-D maze setup only) sequences, two widely-used low-discrepancy Quasi-Monte Carlo methods. Sobol sequences are designed to efficiently cover spaces by minimizing gaps between sample points, making them ideal for complex problems. Halton sequences, on the other hand, generate points using prime number bases, providing effective uniform coverage. However, both Sobol and Halton sequences suffer from correlation issues in higher dimensions [11].
To obtain a statistical measure of performance across repeated runs, we randomize the selection of the MPMC points from the pool of trained point sets (batches can contain 8, 16 or 32 point sets based on the training instance). Similarly for the deterministic Halton and Sobol sequences, we initialize a new sequence where the previous one ended. Hence, when sampling point sets of size $N$, we take the points indexed from $(i-1)\ N+1$ to $i*N$ for the $i$-th run, with $i$ taking integer values from 1 to the number of runs fixed for the experiment.
@@
IV-B Experiments
Figure 3: Visualization of the 2-D maze, SE(3) and UR5 simulation experiments. 2D mazes of increasing difficulty are preseted on the first row: level 1 (left), level 2 (middle) and level 3 (right). The start position is depicted in green and the end goal in orange. The SE(3) cubicle experiment is on the bottom left, with the inverted L shaped object required to navigate the maze, then through the passage to the other side of the board and back to reach the goal configuration (in the bottom right corner). The SE(3) Twistycool setup is in the middle of the bottom row, the red object starting on the left has to rotate its way through the passage in the wall to reach its goal. The 5-D UR5 robot experiment is on the bottom right, it shows the arm reaching into a green box while avoiding contact with the table or box.
IV-B1 2-D maze
The initial experiments consist in using PRM to navigate a series of three 2-D mazes of increasing difficulty depicted in Figure 3. These are designed only to be increasingly difficult to solve for PRM with Uniform sampling before comparing to the other sampling techniques to avoid introducing any bias favoring a specific scheme. The maps are of size 640 by 480 and the agent is assumed to be a disk of radius 6 for collision checking. A single start and end goal is considered and 50 planning attempts are run by level, number of sampled points and sampling technique. The shortest path solver used is $\texttt{A}^{*}$ [39].
IV-B2 OMPL benchmarks
@@
SE(3) rigid body puzzles
The Special Euclidean Group in 3 Dimensions SE(3) is composed of a 3-D translational component and a 3-D rotational component. The latter is often conveniently sampled in quaternion space, with uniform sampling on the 4D unit sphere manifold. In keeping with standard practice, we evaluate samplers’ efficient coverage only on the translational 3-D Euclidean space component. Two such puzzles, Cubicle and Twistycool, are benchmarked from the OMPL. For the latter example, we have reduced the default bounds to allow this harder problem to be solved with the same list of $n$ values used throughout. Both examples can be visualized in Figure 3
$d$-D hypercube corridor
On the $d$-dimensional hypercube, the valid region is defined such that there exists an index $k$ with the following constraints: for all dimensions $i<k$, the $i$-th coordinate must be less than or equal to a threshold edge width $\lambda$, and for all dimensions $i>k$, it must be greater than or equal to $1-\lambda$. This results in a valid subspace that resembles narrow passageways along the hypercube edges, leading from one corner of the hypercube to the diagonally opposite corner. For dimension $d\in\{2,3,10\}$, the value of the edge width is tuned to ensure the point sets available are of sufficient size to establish a meaningful comparison between samplers (respectively $\lambda\in\{0.1,0.2,0.37\}$).
$10$-D kinematic chain
The kinematic chain experiment features a robot with multiple links, operating in a 10-dimensional configuration space. The robot must navigate through an environment with obstacles represented as line segments. Initially, the robot’s first link is at zero radians, and subsequent links are arranged with a specific angular offset. The goal is to move the robot from this starting position to a target configuration where the first link aligns nearly with a desired orientation, while avoiding collisions with the environment and itself.
IV-B3 UR5 robot arm
@@
V RESULTS
V-A 2-D maze
The performance of the different sampling techniques for all three levels of the 2-D mazes are presented in Figure 4. On all three levels and for all fixed number of points, MPMC offers the best value per sampled point set size on all but two instances (outperformed by Halton by 2% on level 1 with $N=32$ and by Sobol by 8% on level 3 with $N=128$). Indeed, MPMC’s efficiency superiority is all the more established on the harder level, where it can reach close to 90% with only 256 points sampled, a feat that takes Halton and Sobol twice the number of points to reach, and a performance level Uniform sampling fails to reach even with 4 times that number. In fact, Uniform sampling is substantially less efficient as it requires sampling at least 4 times the number of points of low discrepancy methods to reach the 90% mark on all levels. Elsewhere, Sobol and Halton sequences seem to offer comparable performance overall, with relative ranking seemingly contingent on the level and number of points sampled. They often fail to meet the performance of MPMC with significant margins (e.g. 34% success for Halton on level 1 with $N=64$ against MPMC’s 54%, under 64% for both Halton and Sobol on level 3 with $N=256$ to MPMC’s 88%).
Figure 4: Success rates of PRM on the 3 levels of 2-D maze motion planning versus the number of points sampled. Results for each level are grouped by color and the data by sampler is identified by marker and line styles.
V-B OMPL and UR5 benchmarks
The benchmarking success rates are provided in Table I, along with the average number of valid points sampled by each scheme and its standard deviation.
On the SE(3) Cubicle and Twistycool experiments, MPMC globally achieves the best performance, although with somewhat marginal advantage. This tightness of margins can be linked to the importance of the SO(3) angle quaternion sampling which is done uniformly in all setups. Halton and Uniform schemes appear to be relatively more contingent on the scenario and are harder to clearly rank.
However, MPMC outperforms both Uniform and Halton samplers more consistently and significantly when used on all the samples’ components. Indeed, over the 2 and 3 dimensionsional hypercube examples, MPMC sampling solves around twice as many runs as its best competitor with 128 and 256 point regimes and maintains an advantage in the region of 50% with larger point sets.
The previous observation remains, to a solid extent, valid in higher dimensions. Although relative performance gains are in multiple cases well reduced, MPMC sampling maintains top-performing status. Furthermore, there remains instances where it offers large performance gains. This is evident on the 10-D kinematic chain test, where MPMC reaches 24% success rate with 256 samples, compared to only 12% for Uniform and 14% for Halton. Similarly, a significant performance gap is established between MPMC (80%) and its counterparts (Halton 56% and Uniform 62%) with 512 points on the 10-D hypercube benchmark. In the realistic scenario of the UR5 planning task (5-D), MPMC consistently surpasses Uniform and Halton sampling (3 to 8-fold higher success rates). We also demonstrate that our approach is readily deployed to real hardware (Figure 5).
@@ V-B OMPL and UR5 benchmarks
TABLE I: Benchmark results on the OMPL suite and UR5 robot arm comparing performance in terms of success rate (SR in %) versus $N$ the number of points sampled $\{128,256,512,1024\}$. Also provided are the mean and standard deviation of the number of valid milestones in each experiment (denoted ${\lvert V\rvert})$

### v2

@@ Abstract
Keywords: Graph Neural Networks, Discrepancy Theory, Motion Planning
1 Introduction
@@
Instead of refining motion planning algorithms, our work targets the fundamental building block underlying these approaches—the core sampling process itself. We introduce a novel sampling strategy based on Message-Passing Monte Carlo (MPMC) [11], a graph neural network architecture trained to generate low-discrepancy point sets on unit hypercubes of arbitrary dimension. We expand the generic MPMC algorithm with the introduction of a novel training objective tailored to high-dimensional spaces, ensuring that generated points are optimally distributed and scalable.
This leads to a potent unbiased state sampling technique, that can be seamlessly integrated into any sampling-based planner, and that requires neither conditioning on the workspace description, nor a steer function, nor past examples, nor start and goal information. This approach provides strong theoretical guarantees and it also outperforms traditional techniques across various benchmarks, in environments of varying complexity and dimensionality. Our contributions are stated as follows:
@@
Our MPMC-based sampling strategy offers a powerful and versatile alternative to traditional sampling methods. By targeting the sampling process directly, we open up new possibilities for improving the efficiency of a wide array of motion planning algorithms in robotic systems.
2 Related work
2.1 Learning for sampling-based planners
@@
2.2 Low-discrepancy constructions
Over the past century, numerous constructions of low-discrepancy point sets and sequences have been proposed. Most constructions are deeply rooted in number theory and abstract algebra. A widely used building block of low-discrepancy constructions is the one-dimensional van der Corput sequence [18] in base $b$, which is generalized to a higher dimensional setting via the Halton sequence  [19]. Each of the $d$ coordinates in a Halton sequence corresponds to a distinct van der Corput sequence in base $b$, with the bases selected to be co-prime. Faure sequences, as introduced in [20], offer a similar construction to Halton sequences but incorporate permutations of the digits in base $b$.
@@ 2.2 Low-discrepancy constructions
Several approaches have recently emerged targeting the construction of points for fixed dimension and number of points. In [28], new low-discrepancy point sets were proposed by optimizing permutations applied to the Halton sequence. Another approach, known as subset selection, was introduced to select $k<N$ points from an $N$-element set that minimize the discrepancy. An exact algorithm for this selection was presented in [29], while a swap-based heuristic approach was employed in [30]. Additionally, [31] proposed a non-linear programming method to generate point sets with optimal star-discrepancy for fixed dimension and number of points. However, this approach faces significant computational challenges, limiting its practical application to finding optimal sets for only up to $21$ points in two dimensions and $8$ points in three dimensions. MPMC significantly differs from all previous approches by its explicit use of machine learning. Moreover, it has been shown in [11] that MPMC generates point sets with significantly better distributional properties compared to any previous method, reaching optimal or near-optimal discrepancy.
3 Methods
@@
3.1 Message-Passing Monte Carlo Sampling
Message-Passing Monte Carlo (MPMC) [11] leverages Graph Neural Networks (GNNs) to generate point sets that cover the space in a uniform manner (for an extension to non-uniform distributions see [32]). This uniformity can be assessed through measures of irregularity termed discrepancy. While there exist a plethora of different uniformity measures, we focus on the ${\mathcal{L}}_{p}$-discrepancy here. That is, given a set of points $\{{\bf X}_{i}\}_{i=1}^{N}$ in the unit hypercube $0,1]^{d}$ and $p\geq 1$, the ${\mathcal{L}}_{p}$-discrepancy is defined as,
${\mathcal{L}}_{p}^{p}(\{{\bf X}_{i}\}_{i=1}^{N})=\int_{[0,1]^{d}}\left|\frac{\#(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,{\bf x}))}{N}-\mu([0,{\bf x}))\right|^{p}d{\bf x},$ (1)
where $(\{{\bf X}_{i}\}_{i=1}^{N}\cap[0,{\bf x}))$ counts how many points of $\{{\bf X}_{i}\}_{i=1}^{N}$ fall inside the box $0,{\bf x})=\prod_{i=1}^{d}[0,x_{i})$ for ${\bf x}=(x_{1},\ldots,x_{d})\in[0,1]^{d}$, and $\mu(\cdot)$ denotes the usual Lebesgue measure. A differentiable closed-form solution to the high-dimensional integral in (1) for the case of $p=2$, known as Warnock’s formula [33], can be computed in $\mathcal{O}(N^{2}d)$. Consequently, MPMC utilizes the ${\mathcal{L}}_{2}$-discrepancy as the training loss to generate point sets with low-discrepancy. The MPMC model transforming random input points into low-discrepancy sets via deep GNNs is depicted in Fig. 1.
3.1.1 Fast MPMC training in higher dimensions
As described in [11], the ${\mathcal{L}}_{2}$-discrepancy fails to distinguish random from highly uniform points in high dimensions for moderate amount of points. Therefore, it was suggested in [11] to minimize the Hickernell ${\mathcal{L}}_{2}$-discrepancy $D_{H,2}$ [34] in this case, that sums over all ${\mathcal{L}}_{2}$-discrepancies of $k$-dimensional projections, with $1\leq k\leq d$. A straightforward implementation of $D_{H,2}$ has a computational complexity of $\mathcal{O}((2^{d}-1)N^{2}d)$, which would involve evaluating the $\mathcal{L}_{2}$-discrepancy for billions of projections for $d\geq 30$. To mitigate this issue, the authors in [11] limit the summation to smaller, randomly selected subsets of projections. In contrast, in this work we leverage the closed-form solution of the Hickernell ${\mathcal{L}}_{2}$-discrepancy from [35] that can be computed again in $\mathcal{O}(N^{2}d)$ instead of $\mathcal{O}((2^{d}-1)N^{2}d)$ [36]. More concretely,
$=\left(\frac{4}{3}\right)^{d}-\frac{2}{N}\sum_{i=1}^{N}\prod_{k=1}^{d}\left(\frac{3}{2}-\frac{{\bf X}_{i,k}^{2}}{2}\right)+\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\prod_{k=1}^{d}[2-\max({\bf X}_{i,k},{\bf X}_{j,k})],$
where ${\bf X}_{i,k}$ is the $k$-th entry of ${\bf X}_{i}$, and $\{{\bf X}^{(s)}_{i}\}_{i=1}^{N}$ is the projection of $\{{\bf X}_{i}\}_{i=1}^{N}$ onto $0,1]^{|s|}$. This enables very fast training in higher dimensions, minimizing the discrepancy over all projections.
3.2 Motion planning via Probabilistic Road Maps
Algorithm 1 ps-PRM Algorithm
@@ 3.2 Motion planning via Probabilistic Road Maps
[…]
@@
The algorithm begins by sampling $N$ points from the entire space to form the set $S$. Next, we prune out milestones from the initial set $S$ that do not fall in free space, keeping only valid samples $S_{v}$, which are combined with the start and goal $\{{\bf X}_{\text{start}},{\bf X}_{\text{goal}}\}$ to obtain $V$. For each node $v$ in $V$, the algorithm identifies nearby nodes within a radius $r_{N}$, with the set of such points denoted $P_{\text{near}}$. For each neighbor $p$ in $P_{\text{near}}$, the algorithm checks for collision-free paths between $v$ and $p$. If the path is free of collisions, a bidirectional edge is added between $v$ and $p$ to the edge set $E$. Finally, the algorithm attempts to find a shortest path from ${\bf X}_{\text{start}}$ to ${\bf X}_{\text{goal}}$ and returns the path if one exists; otherwise, it indicates failure.
3.3 Theoretical guarantee
Next, we outline a theoretical justification of minimizing discrepancy as a means to improving efficiency of sampling-based motion planning. To this end, we introduce another uniformity measure known as dispersion, which is commonly used in assessing the efficiency of sampling-based motion planning. Given a point set $\{{\bf X}^{s}_{i}\}_{i=1}^{N}$ in $0,1]^{d}$ and $p\geq 1$, the $l_{p}$-dispersion is then defined as,
@@ 3.3 Theoretical guarantee
[…]
@@
for some $\alpha>1$ and $r_{N}$ satisfying conditions corresponding to the path $\delta$-clearance, one can guarantee that the cost of the returned path is within a factor of $\frac{1}{\alpha-1}$ of the optimal $\delta$-clear path. Clearly, a low-discrepancy induces a guarantee of the resulting path to be close to the optimal $\delta$-clear path.
Note.
[…]
4 Experimental Evaluation
We evaluate our motion planning solution in simulation, using a variety of environments ranging from 2D mazes to higher dimensional spaces and in a physical setting with a UR5 robot arm. We assess performance according to the number of sampled points and success rate in finding a path.
4.1 Sampling methods
@@
The performance of our method is evaluated against Uniform sampling and also against Halton [39] and Sobol [40] (for the 2-D maze setup only) sequences, two widely-used low-discrepancy Quasi-Monte Carlo methods (see Appendix A.1 for 2D-visualization). Sobol sequences are designed to efficiently cover spaces by minimizing gaps between sample points, making them ideal for complex problems. Halton sequences generate points using prime number bases, providing effective uniform coverage. However, both approaches suffer from correlation issues in higher dimensions [11].
To obtain a statistical measure of performance across repeated runs, we randomize the selection of the MPMC points from the pool of trained point sets (batches can contain 8, 16 or 32 point sets based on the training instance). Similarly for the deterministic Halton and Sobol sequences, we initialize a new sequence where the previous one ended. Hence, when sampling point sets of size $N$, we take the points indexed from (i1)N+1(i-1)\ N+1 to iNi*N for the $i$-th run, with $i$ taking integer values from 1 to the number of runs fixed for the experiment.
@@
4.2 Experiments
Figure 3: Experiments of increasing task complexity: 2D mazes (top) with start (green) and goal (orange) for basic navigation, followed by SE(3) (bottom left/center) for complex object rotation/translation, and 5-D UR5 (bottom right) for constrained robotic manipulation.
2-D maze.
The goal is to navigate a series of three 2-D mazes depicted in Figure 3. These are designed only to be increasingly difficult to solve for PRM with Uniform sampling before comparing to the other sampling techniques to avoid introducing any bias favoring a specific scheme. The maps are of size 640 by 480 and the agent is assumed to be a disk of radius 6 for collision checking. A single start and end goal is considered and 50 planning attempts are run by level, number of sampled points and sampling technique. The shortest path solver used is $\texttt{A}^{*}$ [41].
OMPL benchmarks
@@ 4.2 Experiments
1. SE(3) rigid body puzzles. The Special Euclidean Group in 3 Dimensions SE(3) is composed of a 3-D translational component and a 3-D rotational component. The latter is often conveniently sampled in quaternion space, with uniform sampling on the 4D unit sphere manifold. In keeping with standard practice, we evaluate samplers’ efficient coverage only on the translational 3-D Euclidean space component (discussion on this point provided in Appendix A.4. Two puzzles, Cubicle and Twistycool, are benchmarked from the OMPL. For the latter, we reduce the bounds to allow this harder problem to be solved with the same $N$ values. Both examples can be visualized in Figure 3
2. $d$-D hypercube corridor. On the $d$-dimensional hypercube, the valid region is defined such that there exists an index $k$ with the following constraints: for all dimensions $i<k$, the $i$-th coordinate must be less than or equal to a threshold edge width $\lambda$, and for all dimensions $i>k$, it must be greater than or equal to $1-\lambda$. This results in a valid subspace that resembles narrow passageways along the hypercube edges, leading from one corner of the hypercube to the diagonally opposite corner. For dimension $d\in\{2,3,10\}$, the value of the edge width is tuned to ensure a meaningful comparison between samplers (respectively $\lambda\in\{0.1,0.2,0.37\}$).
3. $10$-D kinematic chain. The experiment features a robot with multiple links in a 10-dimensional configuration space. The robot must navigate through an environment with obstacles represented as line segments. Initially, the robot’s first link is at zero radians, and subsequent links are arranged with a specific angular offset. The goal is to reach a target configuration where the first link aligns nearly with a desired orientation, while avoiding collisions with the environment and itself.
UR5 robot arm
@@
5 Results
5.1 2-D maze
Figure 5: Success rates on the 3 levels of 2-D mazes versus the number of points sampled. Per-level results are grouped by color and the data by sampler is identified by marker and line styles.
The performance of the different sampling techniques for all three levels of the 2-D mazes are presented in Figure 5. On all three levels and for all fixed number of points, MPMC offers the best value per sampled point set size on all but two instances (outperformed by Halton by 2% on level 1 with $N=32$ and by Sobol by 8% on level 3 with $N=128$). Indeed, MPMC’s efficiency superiority is all the more established on the harder level, where it can reach close to 90% with only 256 points sampled, a feat that takes Halton and Sobol twice the number of points to reach, and a performance level Uniform sampling fails to reach even with 4 times that number.
In fact, Uniform sampling is substantially less efficient as it requires sampling at least 4 times the number of points of low-discrepancy methods to reach the 90% mark on all levels. Elsewhere, Sobol and Halton sequences seem to offer comparable performance overall, with relative ranking seemingly contingent on the level and number of points sampled. They often fail to meet the performance of MPMC with significant margins (e.g. 34% success for Halton on level 1 with $N=64$ against MPMC’s 54%, under 64% for both Halton and Sobol on level 3 with $N=256$ to MPMC’s 88%).
5.2 OMPL and UR5 benchmarks
The benchmarking success rates are provided in Table 1, along with the average number of valid points sampled by each scheme and its standard deviation.
On the SE(3) Cubicle and Twistycool experiments, MPMC globally achieves the best performance, although with a marginal advantage. This tightness of margins can be linked to the importance of the SO(3) angle quaternion sampling which is done uniformly in all setups. Halton and Uniform schemes appear to be relatively more contingent on the scenario and are harder to clearly rank.
However, MPMC outperforms both Uniform and Halton samplers more consistently and significantly when used on all the samples’ components. Indeed, over the 2 and 3 dimensional hypercube examples, MPMC sampling solves around twice as many runs as its best competitor with 128 and 256 point regimes and maintains an advantage in the region of 50% with larger point sets.
The previous observation remains, to a solid extent, valid in higher dimensions. Although relative performance gains are in multiple cases well reduced, MPMC sampling maintains top-performing status. Furthermore, there remains instances where it offers large performance gains.
This is evident on the 10-D kinematic chain test, where MPMC reaches 24% success rate with 256 samples, compared to only 12% for Uniform and 14% for Halton. Similarly, a significant gap is established between MPMC (80%) and its counterparts (Halton 56% and Uniform 62%) with 512 points on the 10-D hypercube benchmark. In the realistic scenario of the UR5 planning task (5-D), MPMC consistently surpasses Uniform and Halton sampling (3 to 8-fold higher success rates). We also demonstrate that our approach is readily deployed to real hardware (Figure 4).
[…]